# Longhorn: State Space Models are Amortized Online Learners

### v4

@@ Abstract
The most fundamental capability of modern AI methods such as Large Language Models (LLMs) is the ability to predict the next token in a long sequence of tokens, known as “sequence modeling.” Although the Transformers model is the current dominant approach to sequence modeling, its quadratic computational cost with respect to sequence length is a significant drawback. State-space models (SSMs) offer a promising alternative due to their linear decoding efficiency and high parallelizability during training. However, existing SSMs often rely on seemingly ad hoc linear recurrence designs. In this work, we explore SSM design through the lens of online learning, conceptualizing SSMs as meta-modules for specific online learning problems. This approach links SSM design to formulating precise online learning objectives, with state transition rules derived from optimizing these objectives. Based on this insight, we introduce a novel deep SSM architecture based on the implicit update for optimizing an online regression objective. Our experimental results show that our models outperform state-of-the-art SSMs, including the Mamba model, on standard sequence modeling benchmarks and language modeling tasks. [^1] Code is available at https://github.com/Cranial-XIX/longhorn.git.
Figure 1: (left) The average perplexity across eight downstream datasets for GLA, Mamba and Longhorn (1.3B model) over seen tokens on SlimPajama. Longhorn leads to a 1.8x speed up in sampling efficiency. (right) Longhorn, pretrained with 2048 context length, successfully extrapolates to longer context (up to 16x of the trained context).
@@ 1 Introduction
The Transformer model has become the go-to architecture for sequence modeling in deep learning (Vaswani et al., 2017). However, its utility is constrained by the quadratic growth in training and decoding costs with increasing sequence length. Despite various optimizations such as efficient decoding (Chen et al., 2023; Kuperman & Dyke, 2011), KV-cache compression (DeepSeek-AI & Dai, 2024), and memory efficient implementation (Dao et al., 2022), scaling Transformers for autonomous and continual use with an infinite (or very long) context window remains a challenge.
Recent advancements in linear attention models (Katharopoulos et al., 2020) and state-space models (SSMs)(Gu et al., 2021) have demonstrated their potential. These models are specialized recurrent neural networks capable of efficiently computing outputs in parallel when input tokens are provided simultaneously during training, thus avoiding the inefficiencies of traditional backpropagation through time. During inference, the recurrent form is employed, resulting in linear decoding efficiency. Initially, these models underperformed compared to Transformers. However, recent SSMs (Gu & Dao, 2023; Yang et al., 2023; Peng et al., 2024; De et al., 2024; Beck et al., 2024) have achieved performance parity with Transformers in language modeling tasks. Despite extensive research into various design aspects of SSMs, a guiding principle for designing SSMs remains elusive.
In this work, we propose one potential principle. We observe that one can view SSMs (or any sequence mixing layers) as “meta modules” that compress the history into a memory state which is then used by later layers in the network for sequence modeling. From this perspective:
@@
Figure 2: (left) Most existing sequence models consist of channel and sequence mixing layers. The sequence mixing layers can be viewed as “meta-modules” that compress history into a state $s_{t}$, which is then passed to later layers for sequence modeling. (middle) One can view sequence mixing as addressing an online learning problem, where the state of the SSM follows certain dynamics that optimize an online learning objective. The recurrent update of $s$ can then be derived by solving the online objective in closed form or by conducting a gradient step. (right) Longhorn’s dynamics is derived from solving an online regression problem, where the goal is to recover $x\in\mathbb{R}$ (if $x\in\mathbb{R}^{d}$ we can run $d$ SSMs in parallel) based on a hint $k\in\mathbb{R}^{m}$ from a state vector $s\in\mathbb{R}^{d}$. The closed-form solution to the online regression problem corresponds to the implicit online learning update, where $\Delta_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})$. Please see the details in Section 3.
As a result, we can draw inspiration from online learning and confine the design choices of SSMs to reflect those learning dynamics that solve specific online prediction problems. The hope is that by selecting the right objective, one can save parameters/computation without sacrificing performance. In addition, the proposed online learning perspective might lead to a better understanding of what SSM layers are doing in a large model. Specifically, we view the recurrent update (a.k.a., the state-transition dynamics) of an SSM model as either a gradient descent step or a closed-form solution to an online learning objective. We provide the objectives for some existing SSMs in Table 1.
@@
Based on this insight, we propose a simple yet effective architecture (Longhorn), derived from the implicit closed-form update of an online associative recall problem. The closed-form update naturally leads to a stable recurrent form without a manually designed gating mechanism. So Longhorn does not need a separately parameterized forget gate, which saves parameters when the state size is large. We demonstrate that Longhorn performs comparably to or better than state-of-the-art SSMs like Mamba (Gu & Dao, 2023) on both synthetic and large-scale sequence modeling tasks. In particular, Longhorn outperforms Mamba at the size of 1.3B-parameter when trained on 100B tokens from the SlimPajama dataset (Soboleva et al., 2023). To summarize, our contributions are:
@@
1) Theoretical Framework: We propose a novel framework that views SSM’s recurrent update as solving online learning objectives. As a result, the design of SSMs reduces to the design of the online learning objectives. In particular, we introduce a novel, simple, and effective SSM, named Longhorn, that explicitly solves an online associative recall problem. Longhorn’s recurrent update is obtained by the closed-form solution to the online learning objective. Consequently, Longhorn does not require a separately parameterized forget gate that appears in most existing SSMs.
@@
2) Empirical Results: Longhorn demonstrates better performance than existing SSMs including Mamba, across both synthetic associative recall tasks and the large-scale language modeling task. Moreover, it achieves 1.8x enhancement in sampling efficiency compared to Mamba (See Figure 1 (left)). Longhorn’s training speed is as fast as Mamba, as we only replace the SSM module in the Mamba architecture with Longhorn’s recurrence. So it serves as a drop-in replacement for Mamba. Lastly, Longhorn, trained with 2048 context length can extrapolate to 32K context length at inference time without much perplexity drop (See Figure 1 (right)).
@@
5:     2) Compute the query $q_{t}$, key $k_{t}$ and $\beta_{t}$ (as in objective $||s-s_{t-1}||^{2}+\beta_{t}||s^{\top}k_{t}-x_{t}||^{2}$):
@@
$S_{t}=\big{(}1-\Delta_{t}\otimes k_{t}^{\odot 2}\big{)}\odot S_{t-1}+\big{(}\Delta_{t}\odot x_{t}\big{)}\otimes k_{t},$
where $\Delta_{t}$ is the step size:
$\Delta_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})\in(0,1)^{d}.$
@@ 2 Background
Deep SSMs replace the SA layers with SSM layers. Some variants of SSM models leave the MLP layers unchanged (Sun et al., 2023; Yang et al., 2023; De et al., 2024), while others fuse the SSM layer and the MLP layer into a single unified module (Gu & Dao, 2023). But in both cases, the sequence mixing is done by the SSM module, and the channel mixing is done by the channel-wise MLP. Taking Mamba as an example (Gu & Dao, 2023), a Mamba model consists of a stack of homogeneous modules named Mamba block (the $\Phi(\bm{x})$); we provide a visualization of a single Mamba block in Figure 3 (Gu & Dao, 2023), which consists of an SSM block for sequence mixing (red), and an MLP block for channel mixing (blue).
@@
where $x_{t}$ is the input at time $t$, $S_{t}$ is the model’s state, $A_{t},B_{t}\colon\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are some functions and * is a multiplication operation of choice, such as Hadamard product or matrix product.
@@
A key feature of this design in Equation 1 is that the update of $S_{t}$ is a linear recurrence relation. That is, $S_{t}$ is a linear function of $S_{t-1}$. Crucially, this allows us to express all $S_{t}$ in an explicit form that can be calculated in parallel: when all $\bm{x}=\{x_{t}\}_{t}$ are available as in the training phase, $\{S_{t}\}_{t}$ can be written into
@@
Here $\prod$ denotes the product induced by multiplication operator *. The resulting cumulative product $\overline{A}_{t^{\prime}\to t}$ can be implemented efficiently in parallel with the prefix scan algorithm (Harris et al., 2007), which only requires a sublinear complexity in terms of sequence length (e.g., $\mathcal{O}(\log t)$). From now on, we will abbreviate $A(x_{t})$ and $B(x_{t})$ as $A_{t}$ and $B_{t}$, respectively.
@@
Existing variants of SSMs mainly differs in the design choices of the networks $A_{t},B_{t}$, and the associated operator * in the linear recurrence. A core issue here is that the memory state $S_{t}\in\mathbb{R}^{d\times m}$, designed be $m$ times of the input $x_{t}$ in size, must be as large as possible to maintain sufficient information during recurrence. This makes the architecture design of $A_{t},B_{t}$, both mapping $\mathbb{R}^{d}$ to $\mathbb{R}^{d\times m}$ challenging. A naive linear lay would result in $d\times d\times m$ weights, which is prohibitively large. This makes it necessary to impose certain low-dimensional structures in $A_{t},B_{t}$, which is what differentiates the existing designs of SSM. In the following, we provide some examples in the form of Equation 1 for some existing SSM models.
Example 2.1 (Linear Attention Variants).
Example 2.2 (Mamba (Gu & Dao, 2023)).
$\begin{split}S_{t}&=A_{t}\odot S_{t-1}+B_{t},~{}~{}~{}~{}\text{where}\\ A_{t}&=\exp(A\odot(\Delta(x_{t})\otimes 1)),\qquad B_{t}=(\Delta(x_{t})\odot x_{t})\otimes k(x_{t}).\end{split}$ (3)
where $S_{t}\in\mathbb{R}^{d\times m}$ with $m=16$ by default, $\Delta(x_{t})\in\mathbb{R}^{d}$, $k(x_{t})\in\mathbb{R}^{m}$ linear mappings of $x_{t}$, and $A\in\mathbb{R}^{d\times m}$ is a data independent (not depending on $x_{t}$ trainable weight matrix.
In Mamba, both $A_{t}$ and $B_{t}$ depend on $\Delta(x_{t})$, which represents the step size for the SSM update.
Example 2.3 (Griffin (De et al., 2024)).
In Mamba and the linear attention variants, the outer product serves as a critical rule of lifting vectors to matrices. The recent Griffin architecture abandons the outer product and performs pure elementwise product:
$s_{t}=a(x_{t})\odot s_{t-1}+\sqrt{1-a(x_{t})}\odot i(x_{t})\odot x_{t},$
where $s_{t},a(x_{t}),i(x_{t})$ are all $\mathbb{R}^{d}$. This yields smaller memory states, but in practice, Griffin is combined with local-attention (i.e., the sliding-window self-attention) to strengthen its capability.
Example 2.4 (RWKV (Peng et al., 2023)).
$s_{t}$ $=u_{t}/z_{t}$
Example 2.5 (HGRN2 (Qin et al., 2024a)).
$S_{t}=(1\otimes f(x_{t}))\odot S_{t-1}+i(x_{t})\otimes(1-f(x_{t})).$
@@ 3.1 SSM as Online Learning
We advocate viewing the recurrence of SSM as solving an online learning problem. In online learning, the agent (the algorithm) picks a state $s_{t}$ at time $t$ and then incurs a loss $\ell_{t}(s_{t})$. The goal is to minimize
@@
For instance, consider online linear prediction, where at each step the agent is given an input-label pair $(x_{t},y_{t})$ and $\ell_{t}(s_{t})=\frac{1}{2}||s_{t}^{\top}x_{t}-y_{t}||^{2}$ is the $\ell_{2}$ regression loss, then the problem becomes an online regression problem, and the goal is to successfully predict $y_{t}$ given $x_{t}$ at future time steps.
Online convex optimization (OCP) yields a principled approach to solving Equation 4 when $\ell_{t}$ are convex, by trading-off the “stability” and “plasiticity” (Mermillod et al., 2013). Formally, an online convex programming algorithm updates $s_{t}$ by solving a regularized cost function:
@@
3.2 SSM as Online Learning and The Longhorn Architecture
Under the online learning framework, the design of SSM reduces to the design of $D_{\phi}$ and $\ell_{t}$ in Equation 5. This provides a unified framework for the existing SSM variants. We summarize in Table 1 the online learning interpretation of several existing SSM architectures.
In this work, we explore a highly simplified and natural design called Longhorn guided by the online principle (see the last row of Table 1). In particular, we consider $\{x_{t}\}_{t}$ as the input stream, where $x_{t}\in\mathbb{R}$ is a scalar (as in Mamba, one can consider there are $d$ parallel SSMs if $x_{t}\in\mathbb{R}^{d}$ is a vector). We want to recurrently update hidden states $\{s_{t}\}_{t}$, where $s_{t}\in\mathbb{R}^{m}$ is a vector that aims to summarize the information up to time $t$. We posit the following OCP objective for updating $s_{t}$:
$s_{t}=\operatorname*{arg\,min}_{s\in\mathbb{R}^{m}}||s-s_{t-1}||^{2}+\beta_{t}(x_{t})||s^{\top}k_{t}(x_{t})-x_{t}||^{2}.$ (6)
Here, $k_{t}(x_{t})\in\mathbb{R}^{m}$ is the linear transformation of the input token $x_{t}\in\mathbb{R}$, which represents the key vector in a standard transformer. $\beta_{t}(x_{t})$ controls how much new information about $x_{t}$ we want the model to incorporate. For instance, $\beta_{t}(x_{t})=0$ implies $s_{t}=s_{t-1}$, while a large $\beta_{t}(x_{t})$ implies the model sacrifices stability for plasticity. For simplicity, we will write $k_{t}=k_{t}(x_{t})$ and $\beta_{t}=\beta_{t}(x_{t})$ from now on.
From a high-level, Equation 6 is solving an online prediction problem of learning a weight $s$ to predict $x_{t}$ given $k_{t}$ with a linear model $x_{t}\approx s^{\top}k_{t}$. It is a supervised formulation of the associative memory problem of memorize $(k_{t},x_{t})$ pairs by learning a mapping from $k_{t}$ to $x_{t}$, such that given a key (input) $k_{t}$ the model can retrieve (predict) its corresponding value (label) $x_{t}$.
Fortunately, this simple objective gives a closed-form solution for $s_{t}$, which coincides with the implicit online learning method (Kulis & Bartlett, 2010).

### v5

@@ Abstract
Modern large language models are built on sequence modeling via next-token prediction. While the Transformer remains the dominant architecture for sequence modeling, its quadratic decoding complexity in sequence length poses a major limitation. State-space models (SSMs) present a competitive alternative, offering linear decoding efficiency while maintaining parallelism during training. However, most existing SSMs rely on linear recurrence designs that appear somewhat ad hoc. In this work, we explore SSM design through the lens of online learning, conceptualizing SSMs as meta-modules for specific online learning problems. This approach links SSM design to formulating precise online learning objectives, with state transition rules derived from solving these objectives. Based on this insight, we introduce a novel deep SSM architecture, Longhorn, whose update resembles the closed-form solution for solving the online associative recall problem. Our experimental results show that Longhorn outperforms state-of-the-art SSMs, including the Mamba model, on standard sequence modeling benchmarks, language modeling, and vision tasks. Specifically, Longhorn achieves a 1.8x improvement in sample efficiency compared to Mamba, and can extrapolate over contexts that are up to 16x longer during inference.
Figure 1: (left) The average perplexity on eight downstream datasets for GLA, Mamba, and Longhorn (1.3B model) over seen tokens on SlimPajama. Longhorn leads to a 1.8x speed up in sampling efficiency. (right) Longhorn, pretrained with 2048 context length, extrapolates up to 16x longer context at inference.
@@ 1 Introduction
The Transformer model has become the go-to architecture for sequence modeling in deep learning (Vaswani et al., 2017). However, its utility is constrained by the quadratic growth in training and decoding costs with increasing sequence length. Despite various optimizations such as efficient decoding (e.g., Chen et al., 2023; Kuperman & Dyke, 2011), KV-cache compression (e.g., DeepSeek-AI & Dai, 2024), and memory efficient implementation (e.g., Dao et al., 2022), it remains challenging to scale Transformers for autonomous and continual use with an infinite (or very long) context window.
Recent advances in linear attention models (Katharopoulos et al., 2020) and state-space models (SSMs) (Gu et al., 2021) have demonstrated their potential. These models are specialized recurrent neural networks capable of efficiently computing outputs in parallel when input tokens are provided simultaneously during training, thus avoiding the inefficiencies of traditional backpropagation through time. During inference, the recurrent form is employed, resulting in linear decoding efficiency. Initially, these models underperformed compared to Transformers. However, recent SSMs (e.g., Gu & Dao, 2023; Yang et al., 2023; Peng et al., 2024; De et al., 2024; Beck et al., 2024) have achieved performance parity with Transformers in language modeling tasks. Despite extensive research into various design aspects of SSMs, a guiding principle for designing SSMs remains elusive.
In this work, we propose one potential principle. We observe that one can view SSMs (or any sequence mixing layers) as “meta modules” that compress the history online into a memory state which is then used by later layers in the network for sequence modeling. From this perspective:
@@
Figure 2: (left) Most existing sequence models consist of channel and sequence mixing layers. The sequence mixing layers can be viewed as “meta-modules” that compress history into a state $S_{t}$, which is then passed to later layers for sequence modeling. (middle) Sequence mixing can be seen as an online learning problem, where the SSM state $S_{t}$ optimizes an online objective. The recurrent update of $S_{t}$ is derived either by solving this objective in closed form or via a proximal update. (right) Longhorn’s update solves online associative recall, where the goal is to recover $x\in\mathbb{R}^{d}$ based on a hint $k\in\mathbb{R}^{m}$ from a state matrix $S\in\mathbb{R}^{d\times m}$. Longhorn’s update corresponds to the implicit online learning solution, where $A_{t}=1_{d\times m}-\varepsilon_{t}\otimes k^{\otimes 2}$ and $B_{t}=(\varepsilon\odot x_{t})\otimes k_{t}$, and $\varepsilon_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})$. See the details in Section 3 and Algorithm 1.
As a result, we can draw inspiration from online learning and confine the design choices of SSMs to reflect those learning dynamics that solve specific online prediction problems. The aim is that, by optimizing for the right objective, the model can achieve superior performance with fewer parameters or reduced computational costs. Furthermore, this online learning perspective may offer deeper insights into the function of SSM layers in large models. In particular, the recurrent update (i.e., state-transition dynamics) of an SSM can be interpreted as a proximal update step or a closed-form solution to an online learning objective. We outline the corresponding objectives for several existing SSMs in Table 4. One significant advantage of viewing SSMs through the lens of online learning is their ability to adapt post-training during deployment, allowing them to process arbitrarily long data sequences at inference time.
@@
Based on this insight, we propose a simple yet effective architecture (Longhorn), derived from the implicit closed-form update of an online associative recall problem. The closed-form update naturally leads to a stable recurrent form without a manually designed gating mechanism, automatically balancing forgetting and learning. Thus Longhorn does not need a separately parameterized forget gate, which saves parameters when the state size is large. We demonstrate that Longhorn performs comparably to or better than state-of-the-art SSMs like Mamba (Gu & Dao, 2023) on synthetic and large-scale sequence modeling tasks. In particular, Longhorn outperforms Mamba at the size of 1.3B-parameter when trained on 100B tokens from the SlimPajama dataset (Soboleva et al., 2023). To summarize, our contributions are:
@@
1) Theoretical Framework: We propose a novel framework that views SSMs’ recurrent update as solving online learning objectives. As a result, the design of SSMs reduces to the design of the online learning objectives. In particular, we introduce a novel, simple, and effective SSM, named Longhorn, that explicitly solves an online associative recall problem. Longhorn’s recurrent update is obtained by the closed-form solution to the online learning objective. Consequently, Longhorn does not require a separately parameterized forget gate that appears in most existing SSMs.
@@
2) Empirical Results: Longhorn demonstrates better performance than existing SSMs including Mamba, across both synthetic associative recall tasks and the large-scale language modeling task. Moreover, it achieves 1.8x improvement in sample efficiency compared to Mamba (See Figure 1 (left)). Longhorn’s training speed is as fast as Mamba, as we only replace the SSM module in the Mamba architecture with Longhorn’s recurrence. So it serves as a drop-in replacement for Mamba. Lastly, Longhorn, trained with 2048 context length can extrapolate to 32K context length at inference time without much perplexity drop (See Figure 1 (right)).
@@
5:     2) Compute the query $q_{t}$, key $k_{t}$ and $\beta_{t}$:
@@
$S_{t}=\big{(}1_{d\times m}-\varepsilon_{t}\otimes k_{t}^{\odot 2}\big{)}\odot S_{t-1}+\big{(}\varepsilon_{t}\odot x_{t}\big{)}\otimes k_{t},$ $\varepsilon_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})\in(0,1)^{d}.$
@@ 2 Background
Deep SSMs replace the SA layers with SSM layers. Some variants of SSM models leave the MLP layers unchanged (Sun et al., 2023; Yang et al., 2023; De et al., 2024), while others fuse the SSM layer and the MLP layer into a single unified module (Gu & Dao, 2023). But in both cases, the sequence mixing is done by the SSM module, and the channel mixing is done by the channel-wise MLP. Taking Mamba as an example (Gu & Dao, 2023), a Mamba model consists of a stack of homogeneous modules named Mamba block (the $\Phi_{i}(\bm{x})$); we provide a visualization of a single Mamba block in Figure 3 (Gu & Dao, 2023), which consists of an SSM block for sequence mixing (red), and an MLP block for channel mixing (blue).
@@
where $x_{t}$ is the input at time $t$, $S_{t}$ is the model’s state, $A_{t},B_{t}\colon\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are some functions of the input and * is a multiplication operation of choice, such as Hadamard product or matrix product.
@@
A key feature of this design in Equation 1 is that $S_{t}$ has a linear recurrence, i.e., $S_{t}$ is linear in $S_{t-1}$.
@@
Crucially, this allows us to express all $S_{t}$ in an explicit form that can be calculated in parallel: when all $\bm{x}=\{x_{t}\}_{t}$ are available as in the training phase, $\{S_{t}\}_{t}$ can be written into
@@
Here $\prod$ denotes the product induced by multiplication operator *. The resulting cumulative product $\overline{A}_{t^{\prime}\to t}$ can be implemented efficiently in parallel with the prefix scan algorithm (e.g., Harris et al., 2007), which only requires $\mathcal{O}(\log T)$ ($T$ is the sequence length) parallel operations. From now on, we will abbreviate $A(x_{t})$ and $B(x_{t})$ as $A_{t}$ and $B_{t}$, respectively.
@@
Existing variants of SSMs mainly differ in the design choices of the networks $A_{t},B_{t}$, and the associated operator * in the linear recurrence. A core issue here is that the memory state $S_{t}\in\mathbb{R}^{d\times m}$, designed to be $m$ times the input $x_{t}$ in size, must be as large as possible to maintain sufficient information during recurrence. This makes the architecture design of $A_{t},B_{t}$, both mapping $\mathbb{R}^{d}$ to $\mathbb{R}^{d\times m}$ challenging. A naive linear mapping would result in $d\times d\times m$ weights, which is prohibitively large. This makes it necessary to impose certain low-dimensional structures in $A_{t},B_{t}$, which is the main difference from existing designs of SSMs. In Appendix A, we summarize some existing deep SSM models in the form of Equation 1.
@@ 3.1 SSM as Online Learning
We advocate viewing the recurrence of SSM as solving an online learning problem. In online learning, the agent picks a state $s_{t}$ at time $t$ and then incurs a loss $\ell_{t}(s_{t})$. The goal is to minimize
@@
For instance, consider online linear prediction, where at each step the agent is given an input-label pair $(x_{t},y_{t})$ and $\ell_{t}(s_{t})=\frac{1}{2}||s_{t}^{\top}x_{t}-y_{t}||^{2}$ is the $\ell_{2}$ regression loss, then the problem becomes an online regression problem, and the goal is to successfully predict $y_{t}$ given $x_{t}$ at future time steps, with the key feature that the prediction ($s_{t}$) can change with each new data point.
Online convex programming (OCP) (e.g., Zinkevich, 2003) yields a principled approach to solving Equation 3 when $\ell_{t}$ are convex, by trading-off the “stability” and “plasticity” (e.g., Mermillod et al., 2013). Formally, an online convex programming algorithm updates $s_{t}$ by solving a regularized cost function:
@@
3.2 The Longhorn Architecture
Under the online learning framework, the design of an SSM reduces to the design of $D_{\phi}$ and $\ell_{t}$ in Equation 4. This provides a unified framework for the existing SSM variants. We summarize in Table 4 in Appendix B the online learning interpretation of several existing SSM architectures.
In this work, we explore a highly simplified and natural design called Longhorn guided by the online principle (see the last row of Table 4). In particular, we consider $\{(k_{t},x_{t})\}_{t}$ as the input stream, where $k_{t}\in\mathbb{R}^{m}$ and $x_{t}\in\mathbb{R}^{d}$ are the key-value pairs, just as in the Transformer model (Vaswani et al., 2017). In practice, as in Mamba (Gu & Dao, 2023), $k_{t}=W_{k}x_{t}\in\mathbb{R}^{m}$, where $W_{k}\in\mathbb{R}^{m\times d}$, is a linear mapping from $x_{t}$.
We want to recurrently update hidden states $\{S_{t}\}_{t}$, where $S_{t}\in\mathbb{R}^{d\times m}$ is a matrix that summarizes the information up to time $t$. We posit the following OCP objective for updating $S_{t}$:
$S_{t}=\operatorname*{arg\,min}_{S\in\mathbb{R}^{d\times m}}\left\{||S-S_{t-1}||^{2}_{\text{F}}+||Sk_{t}-x_{t}||^{2}_{\mathrm{diag}(\beta_{t})}\right\}.$ (5)
Here, $||\cdot||_{\text{F}}$ denotes the Frobenius norm of a matrix, $\beta_{t}\in\mathbb{R}^{d}$ is a vector controlling how much new information about $x_{t}$ we want the model to incorporate for $S_{t}$. For instance, $\beta_{t,i}=0$ implies $S_{t,i}=S_{t-1,i}$ (i.e., the $i$-th row of $S$ remains unchanged), while a large $\beta_{t,i}$ implies the model empties some part of $S_{i}$ for incorporating $x_{t,i}$.
From a high-level perspective, Equation 5 is solving an online prediction problem of learning a weight matrix $S$ to predict $x_{t}$ given $k_{t}$ with a linear model $x_{t}\approx S^{\top}k_{t}$. It is a supervised formulation of the associative memory problem of memorize $(k_{t},x_{t})$ pairs by learning a mapping from $k_{t}$ to $x_{t}$, such that given a key (input) $k_{t}$ the model can retrieve (predict) its corresponding value (label) $x_{t}$.
The objective in Equation 5 is motivated by the observation that the self-attention layer of the Transformer exhibits a form of online associative recall (often referred to as the induction head property) (Olsson et al., 2022). This capability has been shown to underpin the model’s ability to perform in-context learning (Brown, 2020). To explain the connection, in-context learning refers to the model’s ability, during inference, to generalize from a set of provided ${(k,x)}$ (question-answer) pairs and apply this understanding to a new question. This closely parallels associative recall, where the model retrieves relevant information from past interactions to address new inputs.
Fortunately, this simple objective gives a closed-form solution for $S_{t}$, which coincides with the implicit online learning method (e.g., Kulis & Bartlett, 2010), according to Theorem 6 (We provide the proof in Appendix C):